Happie 是一家新奇玩意儿商店的老板,他正在为自己的商店组织一次宣传活动。利用他收购的一批极其稀有的 Lubaba 玩偶,他设计了一个游戏,玩家打开“盲盒”,试图在用于游戏的若干个完全相同的盒子中,找到隐藏在其中一个盒子里的 Lubaba 玩偶。 游戏规则如下:
  • 玩家首先支付 $5 开始游戏,并打开 nn 个盒子中的一个。
  • 如果打开的盒子是空的,玩家可以额外支付 $3 打开第二个盒子。
  • 如果第二个盒子也是空的,玩家可以再次支付 $2.50 打开第三个盒子。
  • 如果第三个盒子仍然是空的,玩家就输掉了游戏。未被赢取的玩偶将被捐赠出去,然后游戏重置,供下一位玩家游玩。
由于 Lubaba 玩偶被视为稀有的收藏品,可以假设在每局游戏中,玩家都会不断打开盒子,直到他们赢得玩偶,并且他们有能力支付最多 3 次允许的开盒尝试。 设 XX 为表示玩家在一次游戏中打开的空盒子数量的随机变量。

  1. (a)

    证明 P(X=2)=1n\text{P}(X = 2) = \frac{1}{n}。

  2. (b)

    求 XX 的概率分布,结果用 nn 表示。

  3. (c)

    如果 n=10n = 10,求 P(X≤σ)\text{P}(X \le \sigma),其中 σ\sigma 是 XX 的标准差。

  4. (d)

    如果一个 Lubaba 玩偶的成本价是 $10,求 Happie 应该准备的盒子数量 nn,使得他组织这次宣传活动既不盈利也不亏损。

先不看答案,让 AI 诊断你卡在哪里?

它不会直接告诉你答案,而是问你卡在哪一步看看 AI 是怎么陪你一步步弄懂这道题的

完整解法

7(a)

1. Identify the sequence of events that results in exactly two empty boxes being opened.

For X=2X=2, the player opens exactly 2 empty boxes before the game ends. Since players keep opening boxes until they win the doll or reach the maximum of 3 tries, having exactly 2 empty boxes means the first two boxes were empty and the third box contained the doll.
P(X=2)=P(1st empty)×P(2nd empty)×P(3rd is doll)\text{P}(X=2) = \text{P}(\text{1st empty}) \times \text{P}(\text{2nd empty}) \times \text{P}(\text{3rd is doll})

2. Calculate the probability using without-replacement multiplication.

There is 1 doll and n−1n-1 empty boxes initially.
P(X=2)=(n−1n)×(n−2n−1)×(1n−2)\text{P}(X=2) = \left(\frac{n-1}{n}\right) \times \left(\frac{n-2}{n-1}\right) \times \left(\frac{1}{n-2}\right)
Cancelling out the common factors (n−1)(n-1) and (n−2)(n-2):
P(X=2)=1n  (shown)\text{P}(X=2) = \frac{1}{n}\ \text{ (shown)}

答案

1n\frac{1}{n}

7(b)

1. Determine the possible values of $X$ and calculate $\text{P}(X=0)$ and $\text{P}(X=1)$.

The possible values of XX (number of empty boxes opened) are 0,1,2,30, 1, 2, 3.
P(X=0)=P(doll in 1st box)=1n\text{P}(X = 0) = \text{P}(\text{doll in 1st box}) = \frac{1}{n}
P(X=1)=P(1st empty)×P(2nd doll)=n−1n×1n−1=1n\text{P}(X = 1) = \text{P}(\text{1st empty}) \times \text{P}(\text{2nd doll}) = \frac{n-1}{n} \times \frac{1}{n-1} = \frac{1}{n}

2. Calculate $\text{P}(X=3)$ using the property that the sum of all probabilities is 1.

Since the game ends after a maximum of 3 tries, ∑P(X=x)=1\sum \text{P}(X=x) = 1.
Using P(X=2)=1n\text{P}(X=2) = \frac{1}{n} from part (a):
P(X=3)=1−P(X=0)−P(X=1)−P(X=2)\text{P}(X = 3) = 1 - \text{P}(X=0) - \text{P}(X=1) - \text{P}(X=2)
P(X=3)=1−1n−1n−1n=n−3n\text{P}(X = 3) = 1 - \frac{1}{n} - \frac{1}{n} - \frac{1}{n} = \frac{n-3}{n}

3. Present the probability distribution in a table.

x0123P(X=x)1n1n1nn−3n\begin{array}{|c|c|c|c|c|} \hline x & 0 & 1 & 2 & 3 \\ \hline \text{P}(X=x) & \frac{1}{n} & \frac{1}{n} & \frac{1}{n} & \frac{n-3}{n} \\ \hline \end{array}
答案
x0123P(X=x)1n1n1nn−3n\begin{array}{|c|c|c|c|c|} \hline x & 0 & 1 & 2 & 3 \\ \hline \text{P}(X=x) & \frac{1}{n} & \frac{1}{n} & \frac{1}{n} & \frac{n-3}{n} \\ \hline \end{array}

7(c)

1. Substitute $n = 10$ into the probability distribution of $X$.

From part (b), the probability distribution of XX is:
P(X=0)=1n\text{P}(X = 0) = \frac{1}{n}
P(X=1)=1n\text{P}(X = 1) = \frac{1}{n}
P(X=2)=1n\text{P}(X = 2) = \frac{1}{n}
P(X=3)=n−3n\text{P}(X = 3) = \frac{n-3}{n}

When n=10n = 10:
P(X=0)=110\text{P}(X = 0) = \frac{1}{10}
P(X=1)=110\text{P}(X = 1) = \frac{1}{10}
P(X=2)=110\text{P}(X = 2) = \frac{1}{10}
P(X=3)=710\text{P}(X = 3) = \frac{7}{10}

2. Calculate the expected value $\text{E}(X)$ and variance $\text{Var}(X)$.

E(X)=∑x P(X=x)\text{E}(X) = \sum x \,\text{P}(X=x)
E(X)=0(110)+1(110)+2(110)+3(710)=0+1+2+2110=2.4\text{E}(X) = 0\left(\frac{1}{10}\right) + 1\left(\frac{1}{10}\right) + 2\left(\frac{1}{10}\right) + 3\left(\frac{7}{10}\right) = \frac{0 + 1 + 2 + 21}{10} = 2.4

E(X2)=∑x2 P(X=x)\text{E}(X^2) = \sum x^2 \,\text{P}(X=x)
E(X2)=02(110)+12(110)+22(110)+32(710)=0+1+4+6310=6.8\text{E}(X^2) = 0^2\left(\frac{1}{10}\right) + 1^2\left(\frac{1}{10}\right) + 2^2\left(\frac{1}{10}\right) + 3^2\left(\frac{7}{10}\right) = \frac{0 + 1 + 4 + 63}{10} = 6.8

Var(X)=E(X2)−[E(X)]2\text{Var}(X) = \text{E}(X^2) - [\text{E}(X)]^2
Var(X)=6.8−(2.4)2=6.8−5.76=1.04\text{Var}(X) = 6.8 - (2.4)^2 = 6.8 - 5.76 = 1.04

3. Find the standard deviation $\sigma$ and evaluate $\text{P}(X \le \sigma)$.

σ=1.04≈1.0198\sigma = \sqrt{1.04} \approx 1.0198

Since XX takes non-negative integer values:
P(X≤σ)=P(X≤1.0198)\text{P}(X \le \sigma) = \text{P}(X \le 1.0198)
P(X≤σ)=P(X=0)+P(X=1)\text{P}(X \le \sigma) = \text{P}(X = 0) + \text{P}(X = 1)
P(X≤σ)=110+110=210=0.2\text{P}(X \le \sigma) = \frac{1}{10} + \frac{1}{10} = \frac{2}{10} = 0.2

答案

0.20.2 (or 15\frac{1}{5})

7(d)

1. Determine the probabilities and the total revenue collected for each outcome.

From part (b), the probability distribution for the number of empty boxes opened (XX) is:
P(X=0)=1n\text{P}(X=0) = \frac{1}{n}
P(X=1)=1n\text{P}(X=1) = \frac{1}{n}
P(X=2)=1n\text{P}(X=2) = \frac{1}{n}
P(X=3)=n−3n\text{P}(X=3) = \frac{n-3}{n}

Let RR be the revenue collected per game. According to the rules, the revenue for each outcome is:
If X=0X=0 (finds doll in 1st box): R=5R = 5
If X=1X=1 (finds doll in 2nd box): R=5+3=8R = 5 + 3 = 8
If X=2X=2 (finds doll in 3rd box): R=5+3+2.50=10.50R = 5 + 3 + 2.50 = 10.50
If X=3X=3 (3 boxes empty, game lost): R=5+3+2.50=10.50R = 5 + 3 + 2.50 = 10.50

2. Formulate an equation using the expected revenue and the expected cost.

For Happie to make no profit and incur no loss, the expected revenue per game must equal the cost of the doll used up per game (which is given away or donated either way).
E(R)=Cost per game\text{E}(R) = \text{Cost per game}
5(1n)+8(1n)+10.50(1n)+10.50(n−3n)=105\left(\frac{1}{n}\right) + 8\left(\frac{1}{n}\right) + 10.50\left(\frac{1}{n}\right) + 10.50\left(\frac{n-3}{n}\right) = 10

3. Solve for the number of boxes $n$.

5+8+10.50+10.50(n−3)n=10\frac{5 + 8 + 10.50 + 10.50(n-3)}{n} = 10
23.5+10.50n−31.50n=10\frac{23.5 + 10.50n - 31.50}{n} = 10
10.50n−8n=10\frac{10.50n - 8}{n} = 10
10.50n−8=10n10.50n - 8 = 10n
0.50n=80.50n = 8
n=16n = 16

答案

n=16n = 16

AI 分析,可能出错

用 AI 批改,圈出写错的那一步,并自动将你的薄弱点存入错题本,形成你的错因库。

同一知识点的其他题

  1. 1一位业余作曲家正在编排一段由四个音符和接下来的三个节拍组成的序列。有 7 种可能的音符(标记为 A 到 G)和 5 种可能的节拍(标记为 1 到 5)。音符和节拍的顺序在创作中很重要。求随机选择的序列满足以下条件的概率: 第三个节拍的数字大…
  2. 2凯(Kai)和里奥(Leo)两兄弟是定期训练的双项全能(跑步和骑行)运动员。在每一天中,凯骑行的概率为 34\frac{3}{4},跑步的概率为 35\frac{3}{5},两者都进行的概率为 pp。 平均而言,里奥在一周的 7 天中有…
  3. 3色觉障碍(色盲)是指辨别颜色差异的能力降低。色觉正常的人能够将 6 种颜色(红、橙、黄、绿、蓝、紫)看作 6 个色带。 患有色觉障碍的 Deutan 无法区分以下两对颜色: (I) 红色和绿色, (II) 蓝色和紫色。 对于特定深浅的这 6…
  4. 4一对同胞 Abbey 和 Betty 还有其他六个朋友一起去参加书展。他们八个人随机排成一队等候一位著名作家的签名。求满足以下条件的概率: Abbey 排在队列第一位且 Betty 排在队列第六位, Abbey 排在队列第一位或 Betty…
  5. 5已知某种特定疾病的感染率约为万分之一(每 10 000 人中约有 1 人感染)。针对该疾病的一项检测准确率为 99.9%,即检测结果与真实情况相符的概率为 99.9%。求在已知某人检测结果呈阳性的条件下,此人确实患有该疾病的概率。 研究人员…